Gundam 0083 Stardust Memory - Episode 4: Battle In The Searing Desert Page 1

Gundam 0083 Stardust Memory Episode 4: Battle In The Searing Desert (English title: The Lost Soldiers)

Translated By Tonghyun “Vajra” Kim edited by Noel Gamboa (noelg@pnet01.cts.com) OPERATION X Version 1.5 1/23/92

Narration

The Federation’s Torrington Base was reduced to ruins. Experiencing the death of comrades for the first time, Uraki was shocked. New pilots were stationed on Albion. However, this resulted in a conflict between veteran pilot Monsha and Uraki. Uraki finally won the position of Unit 1 pilot, but the mission of recapturing Unit 2 awaits him. With different thoughts in each heart, the Albion is taking off.

Opening song

The mobile suit carrier Albion cruises over the dry savannah of Africa. Suddenly, a warning siren goes off.

Sinapus

The enemy? Ivan, prepare for evasive maneuvers!

Simon

It’s not the enemy.

Sinapus

What?

Simon

It’s left the mobile suit deck. It’s them again.

Burning

Simon, you’re not saying…

Simon

Yes, just as you feared, Lieutenant.

On the hangar, two groups stand apart. One side is made up of blue- jacketed mobile suit pilots, the other one made up of technicians. Between the two groups stands Mora, who gives Monsha a punch.

Mora

You’ve gone too far!

Alpha

You fight well, fat lady. But I won’t stand around if you keep beating up my friend.

Mora

Let’s get things straight. Did you come here to chase after girls?

Alpha

Cut the bullshit! Who made us come to in this remote region of Africa in first place, anyway?

Technicians

Don’t brag because you’re pilots. Yeah! Right!

Alpha

Hey sis, don’t get me wrong. I’ll talk with Monsha. I’ll tell him to touch your hip the next time. Haha!

Mora grabs Alpha’s arm and flips him. The incident is monitored on a viewscreen by the entire bridge crew.

Alpha

Wanna fight? You ugly wench!

Mora

Oh, yeah?

Mora

If you are a man, be a gentleman like the Captain!

Alpha

I didn’t know you prefer old men.

Mora

Don’t screw with me!

Monsha

Lina baby!

Mora

Want another one?

Burning, embarrased by his pilot’s antics, walks off. He mumbles to himself as he opens the bridge exit door.

Burning

Those idiots, when are they gonna grow up?

Sinapus

Burning.

Burning

Sir?

Sinapus

This ship consists of people with different backgrounds who hardly know each other. The source of further discord among the crew will be a real nuisance.

Burning

I’m really sorry. I’ll make sure that they obey the rules.

The bridge door closes behind Burning.

Burning

I’m the one who really wants to run amok! You idiots!

Simon

I’m really getting tired of pilots. They think they are the ’stars’.

Sinapus

Watch your mouth, Sergeant.

Simon

I’m sorry. (snicker)

Notice that in the following closeup of Sinapus’ torso, he wears pilot wings on his uniform.

Sinapus

We haven’t been able to trace Gato for a week. And I think the view of Africa is getting rather old. [lit. And I don’t think the view of Africa is fascinating to us anymore.]

Maurice

Captain! It’s another call from Jaburo. It’s the third time today.

Sinapus

We have to no choice but to respond!

Sinapus

Sorry… relay the message.

Nina’s monologue

Searching for stolen Unit 2, Albion is wandering in Africa. Residual Minovski particles and mirages of savanna are hiding Unit 2, as if mocking us. The nature of Earth is so rich in its expression. Will I be irritated, too if things continue to go on like this? I can understand others’ feelings…

Sinapus

Ms. Purpleton, Nina Purpleton!

Nina

Yes.

Sinapus

What do you think about this case?

Nina

Eh? This case?

Jaburo general (on monitor)

Unit 2 has already been delivered by Anaheim. Don’t hesitate! I don’t care whether it’s destroyed to pieces or not. Don’t pay attention to what that lunarian is saying.

Nina

H - how can you say that?

Sinapus

I’ll do best. We’ll engage in action right away. Over.

Jaburo general

Wait.

Sinapus

Ms. Purpleton, I’m sorry you had to hear that. [lit. Sorry about the rudeness (of my superior).]

Nina

No… It was nothing.

Halidda

The brass doesn’t understand. Why do we have to be the only ones to do the searching?

Sinapus

Lt. Halidda. There are other search teams besides us in Africa. We’re not the only ones.

Basrof

Only those with connections to Vice Admiral Cowen.

Maurice

Those at Jaburo are worried about being struck by the nuke of Unit 2. This has got them all confused.

Sinapus

All of you! We’re in the middle of an operation. Behave yourselves!

Nina notices the Core Fighter coming towards the ship.

Kou

Nothing in Kilimanjaro area. Request permission to land.

Deck

Roger. Preparing to receive. Please standby.

Kou

Roger.

Monsha and company are playing cards in a corner of the hangar.

Monsha

What’s this hand? You want my neck or something?

Alpha

Don’t get so carried away. It was you who started this game anyway, to kill some time.

Monsha

Some kind of friends you are.

P/a system

Port hangar, prepare for Core Fighter landing. All personnel in position.

Chap

I guess Uraki’s back.

Alpha

Heh, with another empty hand.

Monsha

Uraki’s back? OK… come with me. Let’s play for the last time.

The deck prepares to receive the Core Fighter.

Monsha

I bet half my pay that Uraki doesn’t make it.

Alpha

The money you owe me says Uraki does.

Chap

I’m not interested.

Monsha

Do as you please, “Mr. Ensign.”

Monsha pops open a small panel on the wall and inserts a playing card between several cables. The resulting short circuit shuts down the guide lights inside the hangar. Uraki’s fighter slips, turns, and crashes onto a protective barrier.

Monsha

I won! All my debts are cleared!

Alpha

He’s not too good at this…

Mora

Wait, you guys did it! What the hell are you up to?

Monsha

Hmph. You again? I just “educated” Uraki on purpose.

Mora

That was the only Core Fighter we had. What will you do if it’s damaged?

Monsha

Is it? So a whiz kid straight outta school can’t even make a safe return if there are no lights?

Monsha points to where he thought Kou would be. As he turns and looks, he sees Nina staring him down.

Nina

Lieutenant, the position of Gundam Unit 1 pilot has been already assigned to Ensign Uraki.

Monsha

Nina, I was too drunk then. I won’t make same mistake twice.

Nina

I hate people with no self-control. I hate them!

Monsha

Hate? Nina, that’s cruel.

Nina

Let’s go, Ensign Uraki.

Kou looks as Nina walks away. He then looks at Monsha, who’s burning with jealousy. Kou, his face pale, salutes Monsha then makes a hasty retreat.

Monsha

Hey, are you running away? Uraki, wait. Let’s do it again. Damn!

Chap

She’ll hate you even more for that.

Monsha

I know! You don’t have to tell me! I’m sick and tired of this combat alert. Where the hell are those Jion suckers hiding?

Keith comes running from the bridge. He runs smack into Burning, who’s hobbling along on a crutch.

Keith

Excuse me!

Burning

Hey, why are you in such a hurry?

Keith

Biological emergency! [lit. It’s a physiological phenomenon!]

Burning

Be calm. You won’t die. Stay on the left side (of the corridor)!

Keith runs smack into another person in the distance.

Burning

I knew it, I knew it…

Keith helps Aubil pick up the papers which the latter dropped.

Aubil

You’re really helpless.

Keith

I’m sorry but I’m in hurry. Sorry, sorry…

Keith’s bladder approaches critical mass.

Aubil

Go ahead. I appreciate your help.

As soon as Keith leaves, Aubil looks around suspiciously, then leaves. Burning approaches the area where Keith and Aubil were, and notices a piece of paper on the floor.

Somewhere in Africa, an old diamond mine [which resembles a small volcano from the outside] buzzes with activity. As Jion soldiers watch anxiously, the Atomic Gundam descends on a platform. The hatch pops open, and a grinning Gato steps out. The soldiers cheer and applaud their hero. Later, in the base commander’s office, a toast is made.

Gato

For the resurrection of Jion!

Bitter

For the success of Operation Stardust!

Gato

Excellent.

Bitter

It’s the last bottle I saved for special occasions.

Bitter

You’re a hero of the soldiers, are a man worthy of drinking this champagne. (The Champagne will be happy to be drunk by the Lieutenant Commander, the hero of soldiers.)

Gato

Your Excellency is the example of a warrior. You did a remarkable job of maintaining this base for three years.

Bitter

I lost many soldiers… and we have only ten mobile suits left. This base was made using the ruins of a diamond mine, it didn’t take much effort to make it.

Gato

Not at all. Thanks to Your Excellency, I’ll be able to transport the Gundam to space.

Bitter

I believe you’ll do a great service for Jion. My efforts in saving an HLV weren’t in vain.

HLV = Heavy-Lift Launch Vehicle

Gato

Your Excellency, I’m sorry I can’t discuss operation details, despite your kindness…

Bitter

That’s the nature of an operation. The only thing that’s left for me is to wait for the launching, while concealing ourselves from the enemy so that the HLV can rendezvous with the retrieval fleet safely.

Gato

Your Excellency…

Bitter

Everything is for the New Jion.

Eyecatch

Soldier

There he is!

Aubil is spotted by ship security, who open fire.

Simon

Alert all security personnel. Anaheim technician Aubil is escaping in the port deck section.

Sinapus

I’ve never guessed that he was the spy of Jion… (The captain looks at a piece of paper containing classified information, the same piece of paper which Burning found.)

Halidda

Captain, the Core Fighter hatch is opening.

A soldier enters the hangar, and finds several mechanics lying on the floor. The Core Fighter’s afterburners are already glowing white.

Soldier

Stop!

Aubil escapes in the fighter, successfully evading the Albion’s fire.

Sinapus

Good, he’s now on his way. Scott, don’t lose him.

Scott

Yes, sir.

Nina

Captain… I’m really sorry that Jion’s spy was from Anaheim.

Sinapus

You don’t have to apologize. Thanks to him, we’ll be able to locate the enemy base.

Nina

Yes, sir.

Aubil

Brau Angel to Ginbaraido. Answer me! I escaped from Albion.

Inside the base, Gato and Bitter observe as conveyor belts carry the Atomic Gundam into the HLV.

Bitter

It’s an outdated HLV, but it’ll be enough to carry you to the rendezvous point.

Gato

That will be more than enough. How can I thank you for all this?

Bitter

Haha… Well, the next hour’s launching might be the greatest accomplishment of this base.

Gato

Yes, sir.

Vaal

Excuse me, sir.

Bitter

What’s the matter?

Vaal

It’s a message from Brau Angel stating that he escaped from the enemy.

Gato

Aubil from Anaheim?

Bitter

No! Ignore it. If we reply, the enemy will know the location of this base.

Vaal

Yes, sir.

Bitter

Wait. Even if we ignore it, they’ll eventually find this base from his course. We’re in trouble.

Back on the Albion…

Scott

Captain, just as we thought. Aubil tried to communicate with the remnants of Jion.

Sinapus picks up his phone.

Sinapus

Lieutenant Burning, can you hear me? It’s time for action!

The Albion unfolds its mobile suit catapults.

Burning

Roger. Now we can say goodbye to the binoculars.

Burning

Monsha, the search team, you’re the leader!

Monsha

Roger.

Two custom GMs step off from their maintenance platforms.

Burning

The support team will be the Bait and Adel units. Standby.

Bait

What?

Monsha

Lieutenant, then…?

Burning

Monsha, take Uraki and Keith with you.

Monsha

We’re not going sightseeing!

Gundam 0083 Stardust Memory - Episode 4: Battle In The Searing Desert Page 10

Burning

It’s an order!

Nina

Captain! How can a formation like this be possible?

Sinapus

I left all mobile suit related affairs to Burning.

Monsha

Damn! I don’t need two bedwetters !

Keith

Bedwetters?

Kou’s Gundam steps on the elevator that will raise it to the launch catapult above.

Kou

Lt. Monsha, we’ll accompany you with diapers ready.

Monsha

Hmph. Are you kidding me?

Monsha looks towards the bridge and sees Nina, who’s back is facing towards him.

Monsha

Oh! Nina! With this sub-camera…

Monsha activates one of his cameras and zooms in an upward angle towards Nina’s skirt.

Mora

Nina, beware of a Peeping Tom under your feet!

Nina

What are you doing?!

Sinapus

Launch them!

Monsha and Kou are catapulted. Meanwhile, Aubil continues to search for someone or something.

Aubil

Damn, the same view over and over again.

Two Zakus step out from some foliage.

Aubil

How kind! A reception team?

The two Zakus open fire on the Core Fighter.

Aubil

What the hell are you doing? Stop it!

The fighter is destroyed in midair, but a parachute pops out seconds before the explosion. Meanwhile, Monsha’s mobile suit platoon jogs in the direction of the Core Fighter.

Gundam 0083 Stardust Memory - Episode 4: Battle In The Searing Desert Page 11

Monsha

Where are you? Aliens! I’m gonna have lots of fun with you.

Keith

Lieutenant, the correct term is ’spacenoid.’

Monsha

Well excuse me for being ignorant .

Kou

Lieutenant! Is your command always like this?

Monsha

Ah, Uraki! I’m a very persistent person. Don’t get cocky just ’cuz you beat me once.

Kou

Yes, sir!

Kous spots the two Zakus.

Kou

Lieutenant, 1 1/2 O’clock!

Monsha

Oh well. I’m sick and tired of waiting. Shoot ’em til they drop!

Kou

L - lieutenant! We can’t expect to hit the targets like this!

Monsha

Ensign Uraki, this is an order. Do you understand?

Kou

Yes, I do.

The mobile suits open fire.

Jion pilot

E - enemy fire?

The Zakus turn around and attempt to put distance between them and the Federation suits, leaving Aubil behind.

Aubil

Wait! What about me?

Monsha

Look at that! The enemy will guide us to their base.

Back in the mine…

Vaal

Your Excellency, the booster installation has been completed.

Bitter puts on gloves.

Bitter

Good. If things go well, I’ll be able to sink the enemy ship with this. At worst we’ll be able to hold the enemy until the HLV launches.

Vaal

I wish you a good fight.

Bitter

Lt. Vaal. After I leave, you’ll be responsible for this base.

Vaal

Yes, sir. What’s on your mind?

Bitter

After the HLV is launched successfully, surrender to the enemy.

Vaal

Eh? Your Excellency!

Bitter

If we lose all our mobile suits in this fight, the role of this base will be over.

Vaal

We can still fight!

Bitter puts his hand on Vaal’s shoulder.

Bitter

The soldiers really fought well during the last three years. Our spirit will live on in Operation Stardust.

Vaal

Through Commander Gato, you mean?

Bitter

The rest is up to you.

Vaal

Your Excellency!

Gato, in a pilot’s spacesuit, looks at Bitter with sadness from a distance. Bitter enters his commandertype Zaku II and gives orders.

Bitter

Attention all units. We’ll assault the enemy ship shortly. Concentrate only on the ship.

The Jion suits step out of the mine. Meanwhile, Monsha’s platoon continues the chase.

Monsha

You’re not getting away, spacenoid! Not after you made us wait so long.

Kou

Incoming projectile. Left side!

Monsha

What?

The missile lands squarely on the GM’s shield. The platton looks to the left and sees two more Jion suits, their gun barrels flashing.

Monsha

Shit, reinforcements?

Back on the Albion…

Burning

They were ambushed?

Maurice

Yes, by two Doms.

Sinapus

Too bad, it seems that enemy didn’t take Aubil.

Burning

Yes, sir. To hide the location of the base from us, they led Monsha and the others in another direction.

Sinapus

Send out the Bait and Adel units. Considering the enemy’s actions, Gundam Unit 2 must be there. The enemy base is in the direction that Aubil was originally headed for.

Meanwhile, the three Federation suits are trapped in a natural trench, unable to return fire or escape because of heavy enemy fire.

Monsha

What? We’re caught by the enemy’s diversion? The real base is in the ruins of a diamond mine? What the hell is this?

Kou

Wasn’t it too good to be true? We have to join Lt. Bait.

Monsha

Don’t try to preach!

Maurice

P.S. Lt. Burning is really mad.

Monsha

Oh, no! Retreat!

Keith

How?

Kou

We won’t be able to reach the Albion, no matter what.

Monsha

I don’t care!

On the Albion…

Ivan

Captain, we’ve completed the turn towards the enemy base.

Sinapus

Good. Battle stations, everyone. We’ll attack directly with the Albion.

Burning

Captain, that’s…

Sinapus

We have no choice. The enemy’s diversion was to buy them some time. The enemy will take action shortly.

Burning

To transport the Gundam Unit 2?

An explosion rocks the ship.

Burning

W - what was that?

From atop a nearby hill, the six other Jion mobile suits look down on the Albion.

Bitter

Only two mobile suits to guard the ship? How fortunate! Let’s go !

Alpha and Chap bear the brunt of protecting the Albion from the attacking Jion suits.

Alpha

And they say the ones left behind get all the luck!

Burning

Damn! Monsha and Uraki, what the hell are they doing?

The Jion suits close in on the trench. One Zaku lobs a grenade. The area is rocked by constant explosions.

Monsha

Finally…

Monsha

There’s explosions everywhere! Uraki! Breach the enemy barrier!

Keith

You can’t! Do you want to kill Uraki in this concentrated attack?

Monsha

Uraki has a Gundam which is the best machine among ours. Uraki… the duel with me wasn’t a show, was it? Scared? Then come out. I’ll be happy to take your place. If you don’t like that idea, surrender! Are you going to hand Gato that Gundam, too?

Uraki’s ego approaches critical mass.

Kou

Ensign Uraki, charging !

Uraki’s Gundam somersaults out of the trench, rockets into the air and opens fire with both the rocket launcher and the rifle.

Jion pilot

What?

The four Jion mobile suits are trashed! Monsha and Keith step out of the trench. Monsha, his jaw wide open, stares at the result. Then he regains his attitude.

Monsha

Not a bad job for a new kid on the block. We’re going to rescue the Albion. Follow me, Ensign Uraki.

Kou

Yes, sir!

Keith giggles. At this point, the Albion continues its low altitude approach towards the mine while defending itself from the Jion mobile suits, which are firing rockets at the Albion and succesfully evading counterfire.

Burning

Captain, please bring the ship higher!

Sinapus

No! We can’t abandon Bait and Adel.

Chap continues to have no luck with his Gun Cannon.

Chap

I can’t use the cannon when I have to run like this.

Alpha delivers a burst of concentrated fire on a Dom, which explodes.

Alpha

Not bad, since you’re used to this terrain… (but still not good enough.)

Alpha flashes a finger at the destroyed suit, but his moment of glory is shattered when his GM loses a foot and falls. Bitter and Heinz’s mobile suits press on the attack on the Albion.

Bitter

Don’t pay any attention to them. Attack from the side!

Simon

The enemy base is now within the range of the main cannon.

Scott

That’s it! The ruins of a mine.

Burning

What? Fog?

Nina

That’s a HLV!

The HLV begins to ascend.

Sinapus

Open the main cannon. Lt. Basrof, adjust the sight.

Bitter sees the HLV.

Bitter

Carry on, Commander Gato. All units, charge!

One Jion suit explodes, obviously from something other than the Albion’s weapons.

Bitter

What? From the side?

Monsha’s platoon arrives in the nick of time.

Kou

We’re on time, Lt.Monsha.

Monsha

See that? We’re in perfect sync. [“Kore ga aun no kokyuu tte yatsu yo” uses an ancient Japanese idiom of Buddhist origin that obviously has no English equivalent. “We’re in perfect sync” is a rough parallel.]

Kou

Lieutenant!

Monsha’s platoon looks up and sees the escaping HLV.

Basrof

I can’t adjust the sight because of enemy fire!

Sinapus

Just shoot !

To everyone’s shock, the main cannon misses the HLV.

Bitter

We won. Glory to Operation Stardust!

Bitter and Heinz activate the boosters mounted on the back of their suits. Both mobile suits shoot up and arc towards the Albion’s main bridge.

Sinapus

Booster? Shoot them down!

The Albion’s guns focus on the two Jion suits, which are now higher than the ship and are descending towards the main bridge. The Albion’s guns successfully destroy one of the mobile suits.

Bitter

Heinz!

Bitter evades the defensive fire and lands on the Albion, just below the bridge’s windows. He jumps up and aims his gun into the windows!

Bitter

Feel our spirit!

Nina screams! There is a flash of fire, and Bitter’s MS spins in midair, shot by someone on the ground. Nina looks down below.

Nina

Ensign Uraki?!

Kou, gasping, looks up towards the Albion. Up in space, the HLV goes through several rocket stages. Inside, Gato looks at a gem given by Bitter as a memento. Later, Nina sees Vaal holding a white flag, leading the other Jion soldiers in surrender.

Kou

Spacenoids… That’s our enemy?

END